# The Effectiveness of Curvature-Based Rewiring and the Role of Hyperparameters in GNNs Revisited

### v1

@@ Appendix C Datasets
If the dataset connected disconnected components we report the statistics for the largest connected component selected. The homophily index $\mathcal{H}(G)$ defined by [35] is defined as
$\mathcal{H}(G)=\frac{1}{|V|}\sum_{v\in V}\frac{\text{ Number of }v\text{ 's neighbors who have the same label as }v}{\text{ Number of }v\text{ 's neighbors }}\,.$ (11)
Texas Cornell Wisconsin Chameleon Squirrel Cora Citeseer Pubmed
$\mathcal{H}(G)$ 0.06 0.11 0.16 0.25 0.22 0.83 0.72 0.79
Nodes 135 140 184 832 2186 2485 2120 19717
Edges 251 219 1703 12355 65224 5069 3679 44324
Features 1703 1703 1703 2323 2089 1433 3703 500
Classes 5 5 5 5 5 7 6 3
Directed? Yes Yes Yes Yes Yes No No No
Appendix D Maximum mean test accuracy values
Table 4: For each dataset we take the maximum value obtained for the mean test accuracy. These are the parameter sweeps which obtained the maximum value and correspond to the top outlier in Figure 2. Top result per dataset in bold.
Texas Cornell Wisconsin Chameleon Cora Citeseer Pubmed
None $63.48$ $54.39$ $56.70$ $\mathbf{48.28}$ $\mathbf{80.96}$ $71.37$ $\mathbf{77.53}$
$BFc$ $59.26$ $56.11$ $54.08$ $41.16$ $80.93$ $70.27$ $68.29$
$BFc_{3}$ $59.26$ $54.64$ $54.14$ $42.83$ $73.59$ $65.42$ $76.97$
$BFc_{mod}$ $59.26$ $56.96$ $\mathbf{64.08}$ $43.04$ $57.58$ $71.46$ $77.39$
$JLc$ $59.26$ $53.61$ $54.35$ $43.64$ $69.94$ $67.71$ $73.54$
$AFc_{3}$ $\mathbf{63.63}$ $\mathbf{60.50}$ $62.37$ $46.62$ $79.81$ $70.97$ $77.24$
$AFc_{4}$ $62.07$ $54.36$ $58.46$ $34.27$ $79.89$ $\mathbf{71.85}$ $74.97$
Appendix E Hardware specifications

### v2

@@ Appendix C Datasets
If the dataset connected disconnected components we report the statistics for the largest connected component selected. The homophily index $\mathcal{H}(G)$ defined by Pei et al. (2019) is defined as
$\mathcal{H}(G)=\frac{1}{|\mathbb{V}|}\sum_{v\in\mathbb{V}}\frac{\text{ Number of }v\text{ 's neighbors who have the same label as }v}{\text{ Number of }v\text{ 's neighbors }}\,.$ (11)
Texas Cornell Wisconsin Chameleon Squirrel Actor Cora Citeseer Pubmed
$\mathcal{H}(G)$ 0.06 0.11 0.16 0.25 0.24 0.22 0.83 0.72 0.79
Nodes 135 140 184 832 2186 4388 2485 2120 19717
Edges 251 219 1703 12355 65224 21907 5069 3679 44324
Features 1703 1703 1703 2323 2089 931 1433 3703 500
Classes 5 5 5 5 5 5 7 6 3
Directed? Yes Yes Yes Yes Yes Yes No No No
Appendix D Hardware specifications
@@
Appendix E Saturation analysis of distributions
Figure 13: Saturation analysis for datasets Texas and Cornell (a.-d.) Distribution of the obtained mean test accuracy when including $33\%$, $66\%$, and $100\%$ of the total hyperparameters runs. (e.) Evolution of the Wasserstein distance between two subsequent distributions that include $x\%$ of the total hyperparameter iterations. (f., g.) Mean and standard deviation when including $x\%$ of the total iterations.
Figure 14: Saturation analysis for datasets Wisconsin and Chameleon (a.-d.) Distribution of the obtained mean test accuracy when including $33\%$, $66\%$, and $100\%$ of the total hyperparameters runs. (e.) Evolution of the Wasserstein distance between two subsequent distributions that include $x\%$ of the total hyperparameter iterations. (f., g.) Mean and standard deviation when including $x\%$ of the total iterations.
Figure 15: Saturation analysis for datasets Squirrel and Actor (a.-d.) Distribution of the obtained mean test accuracy when including $33\%$, $66\%$, and $100\%$ of the total hyperparameters runs. (e.) Evolution of the Wasserstein distance between two subsequent distributions that include $x\%$ of the total hyperparameter iterations. (f., g.) Mean and standard deviation when including $x\%$ of the total iterations.
Figure 16: Saturation analysis for dataset Pubmed (a.-d.) Distribution of the obtained mean test accuracy when including $33\%$, $66\%$, and $100\%$ of the total hyperparameters runs. (e.) Evolution of the Wasserstein distance between two subsequent distributions that include $x\%$ of the total hyperparameter iterations. (f., g.) Mean and standard deviation when including $x\%$ of the total iterations.
Figure 17: Saturation analysis for datasets MUTAG and PROTEINS (a.-d.) Distribution of the obtained mean test accuracy when including $33\%$, $66\%$, and $100\%$ of the total hyperparameters runs. (e.) Evolution of the Wasserstein distance between two subsequent distributions that include $x\%$ of the total hyperparameter iterations. (f., g.) Mean and standard deviation when including $x\%$ of the total iterations.